Water flows at a depth of 40 cm in a100-cm-diameter storm sewer. Calculate the flow rate and the mass flux if the average velocity is 3m/s.


1
Expert's answer
2021-11-16T14:16:03-0500

Given:

d= 100 cm = 1m

v=1m/s

solution:

"A= \pi/4 (d)^2"

"A= 3.14\/4(1)^2 = 0.7854m^2"

flow rate

"Q=AV= 0.7854\times 3= 2.356 m^3/s"

mass flux

"M_f= Q\rho= 2.356\times1000kg/m^3"

"M_f= 2356 kg\/s"